OpenStudy (snickers):

what are the slope and vertical intercept of the line x+2y=8?

OpenStudy (amistre64):

when x = 0 y=? slope is -a/b...-1/2

OpenStudy (anonymous):

slope=- coefficient ofx/coefficient of y =-1/2 vertical intercept: 2y=8 y=4

OpenStudy (anonymous):

lol, in a much simpler way: let's write down the equation in terms of y first : \[y = 4 - x/2\] m= 1/2 and vertical intercept is the y-intercept when x= 0, substitute 0 in the equation in the place of x and solve for y. ^_^

OpenStudy (anonymous):

m= -1/2*

OpenStudy (snickers):

so my slope is -1/2 and my vertical int would be 4?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yes ^_^